# Quasiparticle gap renormalization driven by internal and external screening in WS2 device

### v1

@@ 1.1 Device preparation
    A bulk WS2 crystal with dimensions on the order of a millimetre was picked up with a piece of scotch tap and stamped on blue tape (nitto tape). In parallel, a polydimethylsiloxane (PDMS) Gel-Pak (PF-40x40-0170-X4) of ($10\times 50$) mm2 was placed on a glass slide and cut into pieces of ($10\times 2$) mm2. The prepared blue tape was used to exfoliate WS2 on the PDMS pieces, which were subsequently inspected with an optical microscope to identify single layers.
@@
1.2 Extraction of graphene carrier density
@@
The carrier density $n_{\mathrm{Gr}}$ in the graphene top contact is extracted as a function of $V_{\mathrm{G}}$ by fitting the Fermi wavevector $k_{\mathrm{F}}$ in $V_{\mathrm{G}}$-dependent ARPES spectra of the Dirac cone as shown in Fig. 6(a). For a subset of $V_{\mathrm{G}}$ values we measure the circular Fermi surface and fit a circular contour with radius $k_{\mathrm{F}}$ as seen in Fig. 6(b). In addition, we extract momentum distribution curves (MDCs) at $E_{\mathrm{F}}$ and fit a function composed of two Lorentzian peaks on a linear background, as shown in Fig. 6(c). The distance between the fitted peaks corresponds to $2k_{\mathrm{F}}$. In the low doping regime, around $V_{\mathrm{G}}=0$ V, it is not possible to resolve two separate peaks at $E_{\mathrm{F}}$, as seen via the green curve in Fig. 6(c). We instead fit MDC cuts of the dispersion at lower energies and linearly extrapolate the branches to $E_{\mathrm{F}}$ where we then obtain $2k_{\mathrm{F}}$. Combining these methods we obtain $n_{\mathrm{Gr}}=k_{\mathrm{F}}^{2}/\pi$ at each value of $V_{\mathrm{G}}$ directly from the ARPES dispersion of graphene. We find that the lowest value of carrier density we can reliably estimate is 0.35$\times$1012 cm-2.
1.3 Analysis of valence band positions and linewidths
@@
1.4 Analysis of conduction band dispersion
@@
We adjust the parameters $a$, $b$, $\mathcal{I}_{0}$ and $E_{\mathrm{min}}$ in order to achieve an optimum fit of the EDC through the CBM at KW between experiment an simulation. A comparison of the measured ARPES spectrum and corresponding optimum simulation is presented via Figs. 9(a)-(b) for graphene-encapsulated WS2 at $V_{\mathrm{G}}=12$ V. The resulting EDCs at KW are shown in Fig. 9(c), demonstrating that an excellent fit can be achieved using our simple model of the intensity. Next, we proceed to vary $E_{\mathrm{min}}$ for our optimum parameters and thus simulate the effect on the EDC at KW of shifting the CBM, as shown in Fig. 9(d). From the EDCs, we extract the EDC peak position $E_{\mathrm{peak}}$, which is compared with $E_{\mathrm{min}}$ for our parameters in Fig. 9(e). If $E_{\mathrm{peak}}$ and $E_{\mathrm{min}}$ were identical, the markers would follow a linear proportional relationship with a slope of 1 (see dashed line in Fig. 9(e)). We observe that as $E_{\mathrm{min}}$ shifts to lower energies the value of $E_{\mathrm{peak}}$ approaches $E_{\mathrm{min}}$, whereas the deviation between the two values reaches 20 meV around $E_{\mathrm{F}}$ and increases towards higher energies. However, for our CBM estimates in Fig. 3(c) of the main paper any such discrepancy between our fitted EDC peak position and the actual position of the CBM would be well within our error bar of the band gap of ±$\pm 0.03$ eV.
@@
1.5 Gate voltage dependence of band gaps
@@
1.6 Band positions at the vacuum-graphene interface
@@
1.7 Model ab-initio $G_{0}W_{0}$ calculations
@@
For the derivation of the minimal lattice many-body model we use a combination of density functional theory (DFT) calculations, Wannier function basis sets, and constrained random phase approximation (cRPA) calculations[43]. We start from a DFT calculation for freestanding and undoped monolayer WS2 using the Vienna Ab Initio Simulation Package (VASP)[44, 45] using a projector augmented planewave (PAW) basis set[46] and a PBE exchange correlation functional[47]. The lattice constant of WS2 is set to 3.186 Å. For simplicity we neglect spin-orbit coupling effects, as these would only affect our results quantitatively, but not the qualitative trends we are interested in. Afterwards we project the Kohn-Sham bands to Mo $d_{z^{2}},d_{xy},d_{x^{2}-y^{2}}$ orbitals using Wannier 90. From this we get the minimal tight-binding model, which captures all details of the full ab initio band structure. With these Wannier functions, we perform cRPA calculations to obtain partially screened Coulomb interaction matrix elements, which take the screening of all other bands (except those with predominant $d_{z^{2}},d_{xy},d_{x^{2}-y^{2}}$ character) into account. The calculations are done in momentum space using $16\times 16\times 1$ $k$/$q$ grids and 192 bands. The resulting Coulomb interaction data points are used in a last step to fit an analytical model for the density-density Coulomb matrix elements. To this end, we diagonalize the cRPA density-density Coulomb matrix elements, resulting in 3 eigenvalues. Similar to Ref. [42], the leading eigenvalue $U_{1}(\mathbf{q})$ is fitted to the following function
@@
Equipped with an analytical Coulomb model, the subsequent real-frequency $G_{0}W_{0}$ calculations are performed on the minimal tight-binding model, while retaining the full frequency and momentum dependence of the self-energy in order to describe the plasmon satellite features. These calculations are performed within the TRIQS [51] and TPRF [52] codebases, using 1200 frequency points in an energy window of $\pm$ 9 eV around the center of the (undoped) gap, in combination with $64\times 64$ $k$/$q$ grids. The doping is adjusted by changing the Fermi level, which affects the bare polarization and thereby the screening of the Coulomb interaction.

### v2

@@ 2 Supporting Information
    A bulk WS2 crystal with dimensions on the order of a millimetre was picked up with a piece of scotch tap and stamped on blue tape (nitto tape). In parallel, a polydimethylsiloxane (PDMS) Gel-Pak (PF-40x40-0170-X4) of (10×5010$\times$ 5010 × 50) mm2 was placed on a glass slide and cut into pieces of (10×210$\times$ 210 × 2) mm2. The prepared blue tape was used to exfoliate WS2 on the PDMS pieces, which were subsequently inspected with an optical microscope to identify single layers.
@@
4 II. Extraction of Graphene Carrier Density
@@
The carrier density $n_{\mathrm{Gr}}$ in the graphene top contact is extracted as a function of $V_{\mathrm{G}}$ by fitting the Fermi wavevector $k_{\mathrm{F}}$ in $V_{\mathrm{G}}$-dependent ARPES spectra of the Dirac cone as shown in Fig. 6(a). For a subset of $V_{\mathrm{G}}$ values we measure the circular Fermi surface and fit a circular contour with radius $k_{\mathrm{F}}$ as seen in Fig. 6(b). In addition, we extract momentum distribution curves (MDCs) at $E_{\mathrm{F}}$ and fit a function composed of two Lorentzian peaks on a linear background, as shown in Fig. 6(c). The distance between the fitted peaks corresponds to $2k_{\mathrm{F}}$. In the low doping regime, around $V_{\mathrm{G}}=0$ V, it is not possible to resolve two separate peaks at $E_{\mathrm{F}}$, as seen via the green curve in Fig. 6(c). We instead fit MDC cuts of the dispersion at lower energies and linearly extrapolate the branches to $E_{\mathrm{F}}$ where we then obtain $2k_{\mathrm{F}}$. Combining these methods we obtain $n_{\mathrm{Gr}}=k_{\mathrm{F}}^{2}/\pi$ at each value of $V_{\mathrm{G}}$ directly from the ARPES dispersion of graphene. We find that the lowest value of carrier density we can reliably estimate is 0.35$\times$1012 cm-2. Note that the extrapolation method contributes to the increased scatter of the points with respect to the linear fits seen for small $V_{\mathrm{G}}$-values in Fig. 1(e) of the main text.
5 III. Analysis of Valence band positions and linewidths
@@
6 IV. Analysis of conduction band dispersion
@@
We adjust the parameters $a$, $b$, $\mathcal{I}_{0}$ and $E_{\mathrm{min}}$ in order to achieve an optimum fit of the EDC through the CBM at KW between experiment and simulation. A comparison of the measured ARPES spectrum and corresponding optimum simulation is presented via Figs. 9(a)-(b) for graphene-encapsulated WS2 at $V_{\mathrm{G}}=12$ V. The resulting EDCs at KW are shown in Fig. 9(c), demonstrating that an excellent fit can be achieved using our simple model of the intensity. Next, we proceed to vary $E_{\mathrm{min}}$ for our optimum parameters and thus simulate the effect on the EDC at KW of shifting the CBM, as shown in Fig. 9(d). From the EDCs, we extract the EDC peak position $E_{\mathrm{peak}}$, which is compared with $E_{\mathrm{min}}$ for our parameters in Fig. 9(e). If $E_{\mathrm{peak}}$ and $E_{\mathrm{min}}$ were identical, the markers would follow a linear proportional relationship with a slope of 1 (see dashed line in Fig. 9(e)). We observe that as $E_{\mathrm{min}}$ shifts to lower energies the value of $E_{\mathrm{peak}}$ approaches $E_{\mathrm{min}}$, whereas the deviation between the two values reaches 20 meV around $E_{\mathrm{F}}$ and increases towards higher energies. However, for our CBM estimates in Fig. 3(c) of the main paper any such discrepancy between our fitted EDC peak position and the actual position of the CBM would be well within our error bar of the band gap of ±0.03$\pm$ 0.03± 0.03 eV.
@@
7 V. Gate voltage dependence of band gaps
@@
8 VI. Band positions at the vacuum-graphene interface
@@
9 VII. Model Ab-Initio $G_{0}W_{0}$ Calculations
@@
For the derivation of the minimal lattice many-body model we use a combination of density functional theory (DFT) calculations, Wannier function basis sets, and constrained random phase approximation (cRPA) calculations[41]. We start from a DFT calculation for freestanding and undoped monolayer WS2 using the Vienna Ab Initio Simulation Package (VASP)[42, 43] using a projector augmented planewave (PAW) basis set[44] and a PBE exchange correlation functional[45]. The lattice constant of WS2 is set to 3.186 Å. For simplicity we neglect spin-orbit coupling effects, as these would only affect our results quantitatively, but not the qualitative trends we are interested in. Afterwards we project the Kohn-Sham bands to Mo $d_{z^{2}},d_{xy},d_{x^{2}-y^{2}}$ orbitals using Wannier 90. From this we get the minimal tight-binding model, which captures all details of the full ab initio band structure. With these Wannier functions, we perform cRPA calculations to obtain partially screened Coulomb interaction matrix elements, which take the screening of all other bands (except those with predominant $d_{z^{2}},d_{xy},d_{x^{2}-y^{2}}$ character) into account. The calculations are done in momentum space using 16×16×116$\times$ 16$\times$ 116 × 16 × 1 $k$/$q$ grids and 192 bands. The resulting Coulomb interaction data points are used in a last step to fit an analytical model for the density-density Coulomb matrix elements. To this end, we diagonalize the cRPA density-density Coulomb matrix elements, resulting in 3 eigenvalues. Similar to Ref. [40], the leading eigenvalue $U_{1}(\mathbf{q})$ is fitted to the following function
@@
Equipped with an analytical Coulomb model, the subsequent real-frequency $G_{0}W_{0}$ calculations are performed on the minimal tight-binding model, while retaining the full frequency and momentum dependence of the self-energy in order to describe the plasmon satellite features. These calculations are performed within the TRIQS [49] and TPRF [50] codebases, using 1200 frequency points in an energy window of $\pm$ 9 eV around the center of the (undoped) gap, in combination with $\times 64$ 64 $k$/$q$ grids. The doping is adjusted by changing the Fermi level, which affects the bare polarization and thereby the screening of the Coulomb interaction.